Suppose you inherited $400,000 and invested it at 8.25% per year. How much could you withdraw at the end of each of the next 20 years?
Answer #1

Using financial calculator BA II Plus - Input details:

#

I/Y = YTM / Frequency =

8.250000

FV = Future Value or Face Value =

$0.00

N = Number of years x frequency =

20

PV = Present Value =

-$400,000.00

CPT > PMT = Payment =

$41,501.75

Formula: PMT = |PV| x R% x (1+R%)^N / ((1+R%)^N - 1)

PMT = 400000*0.0825* (1+0.0825)^20 / ((1+0.0825)^20 - 1)) =

$41,501.75

Payment per year = $41,501.75